OpenStudy (anonymous):

x^3 x^-3 y^-4 xy^-1 z^-2 x^2 y^0 z^-1 (x^-5)^3 x^-1 need help guise and it's all about zero and negative exponents.

OpenStudy (jtvatsim):

Remember that anything raised to the 0 power is equal to 1.

OpenStudy (jtvatsim):

Negative exponents follow this rule: x^-n = 1/(x^n). So, \[x^{-4}=\frac{1}{x^4}\]

OpenStudy (anonymous):

i can't clearly understand it, but what i'm actually asking are the answers i typed above. i don't know the answers of those.

OpenStudy (jtvatsim):

OK, so for your first question: \[ x^3 x^{-3} y^{-4}\] The part that scares you is the negative exponents right? So we convert them with the rule and get: \[x^3\frac{1}{x^3}\frac{1}{y^4}\] Then, when we mulyiply together we get: \[\frac{x^3}{x^3y^4}\] The x's cancel and we are left with: \[\frac{1}{y^4}\]

OpenStudy (anonymous):

how bout the other equations i asked? cus not all of em are the same you know. i just want to know all of the answers so that i could study them.

OpenStudy (jtvatsim):

Well, just follow the pattern. So, for \[xy^{-1} z^{-2}\] We still are freaked out by the negative exponents. So let's deal with one at a time. What can you write \[y^{-1}\] as by the rule?

OpenStudy (anonymous):

can you please give all of the answers directly? i just want to know it all right now, cus i still need to study for it. thanks by the way. i appreciate it so much.

OpenStudy (jtvatsim):

Well, if you still need to study for it, then you should probably do that first before just getting the answers. It will work out better that way. People don't learn to play a sport well just by watching other people play them. You have to play the game and make mistakes to become proficient at it. The same is true for math. Math is learned by doing, not just by watching. The answer will only help you once, the process will allow you to solve any question. :)

OpenStudy (anonymous):

Oh, okay. Thanks.

OpenStudy (jtvatsim):

Good luck! It will make sense soon! :)

OpenStudy (anonymous):

Thank you so much!